# Resilient Routing: Risk-Aware Dynamic Routing in Smart Logistics via Spatiotemporal Graph Learning

### v1

@@ Abstract
With the rapid development of the e-commerce industry, the logistics network is experiencing unprecedented pressure. Traditional static routing strategies are often unable to tolerate traffic congestion and fluctuating retail demand. In this paper, we propose a Risk-Aware Dynamic Routing(RADR) framework which integrates Spatiotemporal Graph Neural Networks (ST-GNN) with combinatorial optimization. We first construct a logistics topology graph by using the discrete GPS data using spatial clustering methods. Subsequently, a hybrid deep learning model combining Graph Convolutional Network (GCN) and Gated Recurrent Unit (GRU) is adopted to extract spatial correlations and temporal dependencies for predicting future congestion risks. These prediction results are then integrated into a dynamic edge weight mechanism to perform path planning. We evaluated the framework on the Smart Logistics Dataset 2024, which contains real-world Internet of Things(IoT) sensor data. The experimental results show that the RADR algorithm significantly enhances the resilience of the supply chain. Particularly in the case study of high congestion scenarios, our method reduces the potential congestion risk exposure by 19.3% while only increasing the transportation distance by 2.1%. This empirical evidence confirms that the proposed data-driven approach can effectively balance delivery efficiency and operational safety.
@@ III-A Graph Topology Construction
Real-world logistics data consists of continuous GPS trajectories that do not naturally conform to the discrete graph structures required by Graph Neural Networks (GNNs). Directly modeling each GPS coordinate as a graph node is computationally infeasible and introduces severe noise. To address this issue, we apply a data-driven spatial discretization strategy based on K-Means clustering.
Specifically, all GPS points in the trajectory dataset are clustered into $N=10$ spatial regions, where each cluster represents a functional logistics zone. The cluster centroids are treated as graph nodes $V$, and edges $E$ are constructed based on spatial proximity between centroids. The edge weight is defined as the inverse Euclidean distance between connected nodes, yielding a static adjacency matrix that encodes the underlying logistics topology.
@@ IV-A Dataset & Settings
To construct a graph-based learning dataset, we apply a standardized preprocessing pipeline. First, raw GPS coordinates are spatially clustered using K-Means to discretize continuous trajectories into N = 10 functional logistics nodes, which serve as vertices in the logistics graph. Edges are established based on spatial proximity between cluster centroids, forming a static graph topology. Node-level features are aggregated over fixed temporal intervals to generate time-indexed graph snapshots.
@@ IV-C Ablation Study: Validating the Architecture
The comparative results, presented in Table I, offer a clear answer. The GRU-only variant suffered a significant performance drop (MSE increased to 0.048), proving that traffic conditions in logistics hubs are not isolated events—they are highly correlated with their neighbors. Similarly, the GCN-only model failed to capture the evolving trends (MSE 0.052). The full RADR framework achieves the lowest error (0.041), confirming that simultaneously modeling spatial topology and temporal evolution is non-negotiable for high-accuracy risk prediction.
@@
RADR w/o Spatial (GRU-only) 0.048 +17.1%
RADR w/o Temporal (GCN-only) 0.052 +26.8%
RADR (Full Framework) 0.041
@@ IV-D Case Study
Static $0\to 5$ 293.30 172.04 -
RADR $0\to 9\to 6\to 5$ 299.55 138.90 Risk $\downarrow$ 19.3%
The quantitative impact of this decision is detailed in Table II. By taking this detour, the system effectively negotiated a trade-off: we accepted a marginal 2.1% increase in physical driving distance to secure a massive 19.3% reduction in congestion risk exposure. In a real-world supply chain, trading a few hundred meters of fuel for a significantly higher probability of on-time arrival is a winning strategy.

### v2

@@ Abstract
With the rapid development of the e-commerce industry, the logistics network is experiencing unprecedented pressure. Traditional static routing strategies are often unable to tolerate traffic congestion and fluctuating retail demand. In this paper, we propose a Risk-Aware Dynamic Routing(RADR) framework which integrates Spatiotemporal Graph Neural Networks (ST-GNN) with combinatorial optimization. We first construct a logistics topology graph by combining spatial clustering with trajectory-induced connectivity learning to capture implicit road network constraints. Subsequently, a hybrid deep learning model combining Graph Convolutional Network (GCN) and Gated Recurrent Unit (GRU) is adopted to extract spatial correlations and temporal dependencies for predicting future congestion risks. These prediction results are then integrated into a dynamic edge weight mechanism to perform path planning. We evaluated the framework on the Smart Logistics Dataset 2024, which contains real-world Internet of Things(IoT) sensor data. The experimental results show that the RADR algorithm significantly enhances the resilience of the supply chain. Particularly in the case study of high congestion scenarios, our method reduces the potential congestion risk exposure by 17.6% while only increasing the transportation distance by 2.1%.
@@ III-A Graph Topology Construction
Real-world logistics trajectories are continuous and noisy, which do not naturally conform to the discrete graph structures required by graph neural networks. We first apply K-Means clustering on all GPS points to discretize the service area into $N=10$ functional zones. Each cluster centroid is treated as a node $v\in V$.
@@
To construct a realistic base graph, we derive connectivity directly from historical GPS trajectories rather than relying solely on geometric proximity. Specifically, trajectory records are grouped by the real Truck_ID, and only temporally consecutive GPS records within each truck trajectory are used to derive inter-zone transitions. Let $C_{ij}$ denote the number of observed transitions from zone $i$ to zone $j$ across all trajectories. A directed adjacency matrix $A\in\mathbb{R}^{N\times N}$ is then obtained by row-normalizing the transition counts:
$\mathbf{A}_{ij}=\frac{C_{ij}}{\sum_{k}C_{ik}+\epsilon}$ (1)
To suppress spurious connections caused by noise or rare movements, edges with $A_{ij}<\tau$ are pruned, where $\tau=0.01$ in our experiments. The resulting trajectory-induced directed graph captures asymmetric mobility patterns between zones and provides a more realistic abstraction of the underlying logistical network. The normalized transition weights $A_{ij}$ are further used as edge weights in the GCN layer.
@@ IV-A Dataset & Settings
To construct a graph-based learning dataset, we apply a standardized preprocessing pipeline. First, raw GPS coordinates are spatially clustered using K-Means to discretize continuous trajectories into $N=10$ functional logistics nodes, which serve as vertices in the logistics graph. Edges are established based on historical trajectory transition frequencies grouped by Truck_ID, forming a directed graph topology that reflects actual traffic flows. Node-level features are aggregated over fixed temporal intervals to generate time-indexed graph snapshots.
@@ IV-C Ablation Study: Validating the Architecture
The comparative results, presented in Table I, offer a clear answer. The GRU-only variant suffered a significant performance drop (MSE increased to 0.048), proving that traffic conditions in logistics hubs are not isolated events—they are highly correlated with their neighbors. Similarly, the GCN-only model failed to capture the evolving trends (MSE 0.050). The full RADR framework achieves the lowest error (0.038), confirming that simultaneously modeling spatial topology and temporal evolution is non-negotiable for high-accuracy risk prediction.
@@
RADR w/o Spatial (GRU-only) 0.048 +26.3%
RADR w/o Temporal (GCN-only) 0.050 +31.6%
RADR (Full Framework) 0.038
@@ IV-D Case Study
Static $0\to 5$ 293.30 159.86 -
RADR $0\to 9\to 6\to 5$ 299.55 131.60 Risk $\downarrow$ 17.6%
The quantitative impact of this decision is detailed in Table II. By taking this detour, the system effectively negotiated a trade-off: we accepted a marginal 2.1% increase in physical driving distance to secure a massive 17.6% reduction in congestion risk exposure. In a real-world supply chain, trading a few hundred meters of fuel for a significantly higher probability of on-time arrival is a winning strategy.